[figure: 9.7 cm]

RELATION OF MAGNETIC AND ELECTRIC PROPERTIES OF

THIN FERROMAGNETIC FILMS

A. Ivaškauskas, R. Kaliasas, V. Kleiza, O. Purvinis, R.A. Žilinskas Kaunas University of Technology, Klaipėdos 3, Panevėžys, LT 5319 Lithuania

Thin conductive films, obtained by condensing vapour on dielectric substrates in the thicknesses in which their electric conductivity and grains magnitude depend on their thickness. Electric properties of such films are anisotropic. The values of their temperature resistance coefficient β and resistance R are β(γ) and R(γ), γ is the angle between the vector of electric field E⃗ and the anisotropic axis. When the electric current flows in these films, the transversal voltage U_{⊥}(I_{||}, γ) appears on the boundaries perpendicular to E⃗. Here I_{||} is current intensity. The axis of light magnetization (ALM) of ferromagnetic substances also is the anisotropic axis of electrical properties. U_{⊥} ≈ U_{⊥_{m}} ×sin(2γ±α), here U_{⊥_{m}} is amplitude, α is initial angle between E⃗ and the anisotropic axis. While the films are remagnetised in a quasistationary way by changing the angle between the vector of magnetic field intensity and ALM axis, the B(H⃗) and U_{⊥}(H⃗) dependencies describe the corresponding longitudinal and transversal hysteresis loops. The magnitudes U_{⊥}(H⃗) of two-layered systems add when the current flows through the ferromagnetic layers in the magnetic field: U_{⊥} = U_{⊥}(E⃗, γ) + ΔU_{⊥}(H⃗, E⃗, φ, γ), U_{||} = U_{||}(I_{||}, R_{⊥}) + ΔU_{||}(H⃗, E⃗, φ, γ). ΔU_{||}(H⃗, E⃗, φ, γ) and ΔU_{⊥}(H⃗, E⃗, φ, γ) are the result of anisotropic magnetoresistance effect. The grain sizes of the layers showing conduction anisotropy meet the conditions D_{x} ≠ D_{y} ≠ D_{z}.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : O. Purvinis

Address for correspondence : Kaunas University of Technology, Klaipėdos 3, Panevėžys, LT 5319 Lithuania

Email address : opurvi@panf.ktu.lt